Using the table at right, the total product and average product when 3 workers are employed are Average Physical Marginal or Output Product Product O A. 37 and 27, respectively 0 B, 40 and 10, respectively. O c. 36 and 12, respectively. OD. 39 and 13, respectively 10 12 15 52
Answer #1

ANSWER: C

39 and 13, respectively

Average Physical Product(APP)= Total Product(TP)/ Output

APP at 2 units of labor=12=TP at 2 units of labor/2

TP at 2 units of labor=12*2=24

Marginal Product(MP) at 3 units of labor=15

Hence, TP at 3 units of labor is 24+15=39

APP=TP at 3 units of labor/Output=39/3=13